Please help me out with this maths.
The sum of the digits of a two-digit number is 12. The tens digit is the square of its units digit. Find the number.

Answer by rothauserc(4718)

x + y = 12, where x is the tens digit and y is the units digit
x = y^2
substitute for x in first equation
y^2 + y - 12 = 0
factor the equation
(y-3)*(y+4) = 0
y = 3 or y = -4
x = 9 or x = 16
*****************************************
we have one solution since we are looking for a two digit number
93